The sum of four consecutive even integers is equal to seven times the first integer? What are these integers?
Answer by ReadingBoosters(3246)

x + x+1 + x+2 + x+3 = 7x
4x + 6 = 7x
4x - 4x + 6 = 7x - 4x
6 = 3x
6/3 = 3x/3
2 = x
integers are 2, 3, 4 and 5